OpenStudy (anonymous):

What is the sum of 75 minutes and 55 minutes? 20 minutes 1 hour, 50 minutes 2 hours 2 hours, 10 minutes

OpenStudy (anonymous):

75 + 55 = ?

OpenStudy (anonymous):

2 hours and 10 minutes

OpenStudy (anonymous):

75+55= 130/60 = 2 hrs and 10 min

OpenStudy (clongoria2):

@kliedako You really aren't suppose to give direct answers.

OpenStudy (anonymous):

^Correct.

OpenStudy (anonymous):

Yes, you aren't supposed to give direct answers, even with explanations. Great job @Clongoria2 :)

OpenStudy (anonymous):

ooops sorry, didn't realize that.

OpenStudy (anonymous):

You have to guide them to the answer.

OpenStudy (anonymous):

75+55=130 2 hours and 10 minutes 60=1 hour 60+60=120 10 left 2 hours and 10 minutes